OpenStudy (anonymous):

**WILL MEDAL** Plzz Check My Answer!!!???? What is the solution to the system of equations represented by the two equations? y=-2x+7 y=1/3x ( 3,2 ) ^^^Am I right?

jimthompson5910 (jim_thompson5910):

Incorrect. You should test any potential solutions in both equations y=-2x+7 2=-2(3)+7 ... plug in (x,y) = (3,2) 2 = -6+7 2 = 1 ... this is false so because (3,2) doesn't work for y=-2x+7, it doesn't work for the whole system

OpenStudy (anonymous):

I used this>> https://www.desmos.com/calculator and they intersect at (3,1)

jimthompson5910 (jim_thompson5910):

y=-2x+7 1=-2(3)+7 ... plug in (x,y) = (3,1) 1 = -6+7 1 = 1 ... this is true ------------------------------------------------------- y = (1/3)x 1 = (1/3)*3 ... plug in (x,y) = (3,1) 1 = 1 ... this is true ------------------------------------------------------- So (3,1) is indeed the solution